# Designing LLM Chains by Adapting Techniques from Crowdsourcing Workflows

### v3

@@ 8.4. LLM workflow design processes are unstructured and design support is needed
As there are many overlapping tactics among workflows, the flexibility and reusability of modules can support the development of future workflows (Josifoski et al., 2023). A repository of “battle-tested” workflow operators could allow sharing of quality parts for reuse in new workflows. Such a platform could include operator-level test suites to assist with validation that would monitor operator output quality in new content areas or with newly released LLMs. The platform design could draw on crowdsourcing to create such test suites. Additionally, corresponding workflow design tools could suggest different options to achieve a goal by collecting multiple approaches for sub-parts of the workflow.
@@ 9. Limitations
Scope. Our work targets LLM chain design by adapting crowdsourcing workflow techniques. However, there are opportunities to look beyond LLMs to multi-modal models as the workers (Gupta and Kembhavi, 2023; Surís et al., 2023; Liu et al., 2022) and beyond crowdsourcing to other task decomposition methods as a source of inspiration (Anya, 2015; Andreas et al., 2016; Mao et al., 2019).
@@ 10. Conclusion
By revisiting the established field of crowdsourcing workflows, we find inspiration and direction for the nascent area of LLM chaining. To facilitate adaptation, we contribute a design space for LLM chain designers that incorporates findings from the crowdsourcing workflow and LLM chaining literatures. This design space covers the objectives supported by and the tactical building blocks of prior workflows and chains. We also surface strategies employed by designers to use tactics to support objectives. Through three case studies, we build an understanding for how crowdsourcing workflows can be adapted for LLM chains.
Finally, we synthesize our findings to discuss tactical design decisions for LLM chains through the context of their comparison with crowdsourcing workflows. We find LLM chains need additional support in subtask design and validation due in part to differences between LLM stochasticity and human diversity. On the other hand, the comparatively low cost and latency of LLM calls allows for faster iteration and for workflow techniques that were studied but never practical with human crowds. We hope these findings guide future work towards more approaches to user guidance and creative tasks, systematic discovery of good practices for tactic design, and artifacts to support the chain design process.
LLM chains present remarkable flexibility, which comes along with both promising capabilities but also substantial complexity. Our work helps LLM chain designers to navigate this vast design space by building on the established field of crowdsourcing.
@@ Acknowledgements.
We thank Michael Bernstein and Joy Kim for providing outputs from the original Soylent and Mechanical Novel papers. We also thank the Allen Institute for Artificial Intelligence for access to Amazon Mechanical Turk resources. We thank OpenAI for offering research credits to the GPT-3.5 API. This research was supported in part by NSF award IIS-1901386, ONR grant N00014-21-1-2707, and the Allen Institute for Artificial Intelligence (AI2).References
@@
- Arora et al. (2022) Simran Arora, Avanika Narayan, Mayee F Chen, Laurel J Orr, Neel Guha, Kush Bhatia, Ines Chami, Frederic Sala, and Christopher Ré. 2022. Ask me anything: A simple strategy for prompting language models. arXiv preprint arXiv:2210.02441 (2022).
- Bae et al. (2022) S Sandra Bae, Clement Zheng, Mary Etta West, Ellen Yi-Luen Do, Samuel Huron, and Danielle Albers Szafir. 2022. Making data tangible: A cross-disciplinary design space for data physicalization. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. 1–18.
@@
- Hata et al. (2017) Kenji Hata, Ranjay Krishna, Li Fei-Fei, and Michael S Bernstein. 2017. A glimpse far into the future: Understanding long-term crowd worker quality. In Proceedings of the 2017 ACM Conference on Computer Supported Cooperative Work and Social Computing. 889–901.
@@
- Huang et al. (2023a) Qing Huang, Yishun Wu, Zhenchang Xing, He Jiang, Yu Cheng, and Huan Jin. 2023a. Adaptive Intellect Unleashed: The Feasibility of Knowledge Transfer in Large Language Models. arXiv preprint arXiv:2308.04788 (2023).
@@
- Madaan et al. (2023a) Aman Madaan, Pranjal Aggarwal, Ankit Anand, Srividya Pranavi Potharaju, Swaroop Mishra, Pei Zhou, Aditya Gupta, Dheeraj Rajagopal, Karthik Kappaganthu, Yiming Yang, Shyam Upadhyay, Mausam, and Manaal Faruqui. 2023a. AutoMix: Automatically Mixing Language Models. https://api.semanticscholar.org/CorpusID:264306222
- Madaan et al. (2023b) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, et al. 2023b. Self-refine: Iterative refinement with self-feedback. arXiv preprint arXiv:2303.17651 (2023).
@@
- Sorensen et al. (2024) Taylor Sorensen, Jared Moore, Jillian Fisher, Mitchell Gordon, Niloofar Mireshghallah, Christopher Michael Rytting, Andre Ye, Liwei Jiang, Ximing Lu, Nouha Dziri, et al. 2024. A Roadmap to Pluralistic Alignment. arXiv preprint arXiv:2402.05070 (2024).
@@
- Wu et al. (2023c) Tongshuang Wu, Haiyi Zhu, Maya Albayrak, Alexis Axon, Amanda Bertsch, Wenxing Deng, Ziqi Ding, Bill Guo, Sireesh Gururaja, Tzu-Sheng Kuo, et al. 2023c. LLMs as Workers in Human-Computational Algorithms? Replicating Crowdsourcing Pipelines with LLMs. arXiv preprint arXiv:2307.10168 (2023).
@@
- Wu et al. (2023b) Yiran Wu, Feiran Jia, Shaokun Zhang, Qingyun Wu, Hangyu Li, Erkang Zhu, Yue Wang, Yin Tat Lee, Richard Peng, and Chi Wang. 2023b. An Empirical Study on Challenging Math Problem Solving with GPT-4. arXiv preprint arXiv:2306.01337 (2023).
- Xie et al. (2023) Yuxi Xie, Kenji Kawaguchi, Yiran Zhao, Xu Zhao, Min-Yen Kan, Junxian He, and Qizhe Xie. 2023. Decomposition enhances reasoning via self-evaluation guided decoding. arXiv preprint arXiv:2305.00633 (2023).
@@
- Yao et al. (2023) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Thomas L Griffiths, Yuan Cao, and Karthik Narasimhan. 2023. Tree of thoughts: Deliberate problem solving with large language models. arXiv preprint arXiv:2305.10601 (2023).
@@
- Zhang et al. (2024) Shaokun Zhang, Jieyu Zhang, Jiale Liu, Linxin Song, Chi Wang, Ranjay Krishna, and Qingyun Wu. 2024. Training Language Model Agents without Modifying Language Models. International Conference on Machine Learning (2024).

### v4

@@ 9.4. LLM workflow design processes are unstructured and design support is needed
As there are many overlapping tactics among workflows, the flexibility and reusability of modules can support the development of future workflows (Josifoski et al., 2023). A repository of “battle-tested” workflow operators could allow for sharing of quality parts for reuse in new workflows. Such a platform could include operator-level test suites to assist with validation that would monitor operator output quality in new content areas or with newly released LLMs. The platform design could draw on crowdsourcing to create such test suites. Additionally, corresponding workflow design tools could suggest different options to achieve a goal by collecting multiple approaches for sub-parts of the workflow.
@@ 10. Limitations
Scope. Our work targets LLM chain design by adapting crowdsourcing workflow techniques. However, there are opportunities to look beyond LLMs to multi-modal models and other tools as the workers (Gupta and Kembhavi, 2023; Surís et al., 2023; Liu et al., 2022) and beyond crowdsourcing to other task decomposition methods as a source of inspiration (Anya, 2015; Andreas et al., 2016; Mao et al., 2019).
@@ 11. Conclusion
By revisiting the established field of crowdsourcing workflows, we find inspiration and direction for the nascent area of LLM chaining. To facilitate adaptation, we contribute a design space for LLM chain designers that incorporates findings from the crowdsourcing workflow and LLM chaining literatures. This design space covers the objectives supported by and the tactical building blocks of prior workflows and chains. We also surface strategies employed by designers to use tactics to support objectives. Through three case studies, we build an understanding of how crowdsourcing workflows can be adapted for LLM chains.
Finally, we synthesize our findings to discuss tactical design decisions for LLM chains through the context of their comparison with crowdsourcing workflows. We find LLM chains need additional support in subtask design and validation due in part to differences between LLM stochasticity and human diversity. On the other hand, the comparatively low cost and latency of LLM calls allow for faster iteration and for workflow techniques that were studied but never practical with human crowds. We hope these findings guide future work towards more approaches to user guidance and subjective tasks, the systematic discovery of good practices for tactic design, and artifacts to support the chain design process.
LLM chains present remarkable flexibility, which comes along with both promising capabilities but also substantial complexity. Our work helps LLM chain designers navigate this vast design space by building on the established field of crowdsourcing.
@@ Acknowledgements.
We thank Michael Bernstein and Joy Kim for providing outputs from the original Soylent and Mechanical Novel papers. Thank you to Clement Zheng for your input on creating corpus tables similar to prior work (Bae et al., 2022b). We also thank the Allen Institute for Artificial Intelligence for access to Amazon Mechanical Turk resources. We thank OpenAI for offering research credits to the GPT-3.5 API. This research was supported in part by NSF award IIS-1901386, ONR grant N00014-21-1-2707, and the Allen Institute for Artificial Intelligence (AI2).References
@@
- Arora et al. (2023) Simran Arora, Avanika Narayan, Mayee F Chen, Laurel Orr, Neel Guha, Kush Bhatia, Ines Chami, and Christopher Re. 2023. Ask Me Anything: A simple strategy for prompting language models. In The Eleventh International Conference on Learning Representations. https://openreview.net/forum?id=bhUPJnS2g0X
- Bae et al. (2022a) S Sandra Bae, Clement Zheng, Mary Etta West, Ellen Yi-Luen Do, Samuel Huron, and Danielle Albers Szafir. 2022a. Making data tangible: A cross-disciplinary design space for data physicalization. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems. 1–18.
- Bae et al. (2022b) S. Sandra Bae, Clement Zheng, Mary Etta West, Ellen Yi-Luen Do, Samuel Huron, and Danielle Albers Szafir. 2022b. Making Data Tangible: A Cross-disciplinary Design Space for Data Physicalization. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (New Orleans, LA, USA) (CHI ’22). Association for Computing Machinery, New York, NY, USA, Article 81, 18 pages. https://doi.org/10.1145/3491102.3501939
@@
- Huang et al. (2023a) Qing Huang, Yishun Wu, Zhenchang Xing, He Jiang, Yu Cheng, and Huan Jin. 2023a. Adaptive Intellect Unleashed: The Feasibility of Knowledge Transfer in Large Language Models. CoRR abs/2308.04788 (2023). https://doi.org/10.48550/ARXIV.2308.04788 arXiv:2308.04788
@@
- Lam et al. (2024) Michelle S Lam, Janice Teoh, James A Landay, Jeffrey Heer, and Michael S Bernstein. 2024. Concept Induction: Analyzing Unstructured Text with High-Level Concepts Using LLooM. In Proceedings of the CHI Conference on Human Factors in Computing Systems. 1–28.
@@
- Li et al. (2024) Guohao Li, Hasan Abed Al Kader Hammoud, Hani Itani, Dmitrii Khizbullin, and Bernard Ghanem. 2024. CAMEL: communicative agents for ”mind” exploration of large language model society. In Proceedings of the 37th International Conference on Neural Information Processing Systems (New Orleans, LA, USA) (NIPS ’23). Curran Associates Inc., Red Hook, NY, USA, Article 2264, 18 pages.
@@
- Madaan et al. (2023) Aman Madaan, Pranjal Aggarwal, Ankit Anand, Srividya Pranavi Potharaju, Swaroop Mishra, Pei Zhou, Aditya Gupta, Dheeraj Rajagopal, Karthik Kappaganthu, Yiming Yang, Shyam Upadhyay, Mausam, and Manaal Faruqui. 2023. AutoMix: Automatically Mixing Language Models. https://api.semanticscholar.org/CorpusID:264306222
- Madaan et al. (2024) Aman Madaan, Niket Tandon, Prakhar Gupta, Skyler Hallinan, Luyu Gao, Sarah Wiegreffe, Uri Alon, Nouha Dziri, Shrimai Prabhumoye, Yiming Yang, et al. 2024. Self-refine: Iterative refinement with self-feedback. Advances in Neural Information Processing Systems 36 (2024).
@@
- Shankar et al. (2024) Shreya Shankar, Aditya G Parameswaran, and Eugene Wu. 2024. DocETL: Agentic Query Rewriting and Evaluation for Complex Document Processing. arXiv preprint arXiv:2410.12189 (2024).
- Shao et al. (2024) Rulin Shao, Jacqueline He, Akari Asai, Weijia Shi, Tim Dettmers, Sewon Min, Luke Zettlemoyer, and Pang Wei Koh. 2024. Scaling Retrieval-Based Language Models with a Trillion-Token Datastore. arXiv preprint arXiv:2407.12854 (2024).
@@
- Simkute et al. (2024) Auste Simkute, Lev Tankelevitch, Viktor Kewenig, Ava Elizabeth Scott, Abigail Sellen, and Sean Rintel. 2024. Ironies of Generative AI: Understanding and mitigating productivity loss in human-AI interactions. arXiv preprint arXiv:2402.11364 (2024).
@@
- Wu et al. (2023b) Tongshuang Wu, Haiyi Zhu, Maya Albayrak, Alexis Axon, Amanda Bertsch, Wenxing Deng, Ziqi Ding, Bill Guo, Sireesh Gururaja, Tzu-Sheng Kuo, et al. 2023b. LLMs as Workers in Human-Computational Algorithms? Replicating Crowdsourcing Pipelines with LLMs. arXiv preprint arXiv:2307.10168 (2023).
@@
- Wu et al. (2024) Yiran Wu, Feiran Jia, Shaokun Zhang, Hangyu Li, Erkang Zhu, Yue Wang, Yin Tat Lee, Richard Peng, Qingyun Wu, and Chi Wang. 2024. MathChat: Converse to Tackle Challenging Math Problems with LLM Agents. In ICLR 2024 Workshop on Large Language Model (LLM) Agents.
- Xie et al. (2023) Yuxi Xie, Kenji Kawaguchi, Yiran Zhao, Xu Zhao, Min-Yen Kan, Junxian He, and Qizhe Xie. 2023. Decomposition Enhances Reasoning via Self-Evaluation Guided Decoding. arXiv:2305.00633 [cs.CL]
@@
- Yao et al. (2024) Shunyu Yao, Dian Yu, Jeffrey Zhao, Izhak Shafran, Tom Griffiths, Yuan Cao, and Karthik Narasimhan. 2024. Tree of thoughts: Deliberate problem solving with large language models. Advances in Neural Information Processing Systems 36 (2024).
@@
Appendix A Paper categorization criteria
To show the design space’s completeness in representing the existing literature, we revisited each paper and categorized it within the design space. We describe the nuances of that process here.
We record the year of publication, such that if we collected a paper on arXiv in 2023 and it was published in a conference in 2024, we record the year as 2024.
We show the quality-cost objective spectrum as four columns, in order to show the increasing and decreasing of quality and cost. A paper may both increase and decrease quality or cost at once (e.g., the output is higher quality, but there are reported quality issues with cascading errors). We count the ability to do a task at all as an increase in quality, and an equal match with baseline capabilities as neither an increase nor a decrease. Another challenge for comparison is that the baseline differs across papers. For instance, some papers compare quality to expert outputs, others to novice crowdworkers, and others to zero-shot LLM outputs. We label changes in quality and cost as they are discussed in the paper, rather than use a consistent cross-paper standard. For instance, if the chain could logically be expected to be more expensive than a single call to an LLM, but this is not explicitly discussed in the paper, we do not mark it as increasing costs. We take this approach to avoid speculation and to show trends along with what effects are considered worth reporting.
We used our discretion in classifying papers as only subjective, if the main focus of the task is highly subjective, such as with creative writing. There may still be objective elements, such as wanting a generated story to be grammatically correct. This subjectivity in the categorization itself is why we express it as a spectrum in the design space.
In the design space text, we mention some automated elements that perform similar functions (e.g., rule based heuristics for evaluate, sourcing outside material for generate, and algorithmic merges). We include these in the text descriptions to reference how workflows may use actors or automated elements to complete tasks. However, we do not label a paper as including a subtask of these types if it is only deterministically automated.
Finally, if there are multiple workflows tested in the paper, we include takeaways and elements from all in the coding.